OpenStudy (anonymous):

find the equation of the tangent to y = e^sinx at the point (0, 1)

myininaya (myininaya):

\[y=e^{\sin(x)}\] You need to find y' at (0,1) to find the slope

myininaya (myininaya):

\[\ln(y)=\ln(e^{\sin(x)})\] Now rewrite the right hand side Differentiate both sides Then multiply y on both sides

OpenStudy (anonymous):

thanx alot myininaya